if the first 2 terms of a geometric sequence are √2 and ³√2, what is the fourth term
Answer by sravanik2005(14)

common ratio=2^(1/3)/2^(1/2) =2^(1/3-1/2)=2^(-1/6)
and first term=2^1/2
so,fourth term=2^(1/2) 2^((-1/6)*3))
=2^(1/2) 2^(-1/2) = 2^0 = 1